### 16 Mar 2026 v1

@@ Abstract
Context. Very low-luminosity objects in nearby star-forming regions have been identified as promising proto-brown dwarf candidates. The study of their multiplicity can shed light on the dominant formation mechanism of these substellar objects.
Aims. We aim at studying the multiplicity of the very low luminosity object IRAM 04191+1522.
Methods. We have obtained 0.89 mm ALMA observations with a very extended configuration, achieving an angular resolution of 004 (6 au at 140 pc). We have complemented our data with new VLA observations, and ALMA archival data at 1.3 mm.
Results. We resolve IRAM04191+1522 into a close binary candidate for the first time. The binary is detected in the ALMA continuum data with a projected separation of 80 mas, or 11 au at a distance of 140 pc. The two sources are oriented in the East-West direction, with the eastern component being brighter and more extended than the western one, which is marginally resolved. The analysis of C18O (2-1) archival data reveals gaseous material in rotation around the binary, presumably from a circumbinary disk with 27 au of radius centered on the faintest ALMA component. A fit of the position-velocity diagram allows us to estimate a total dynamical mass for the system of 5040 . Therefore, we classify IRAM04191 as a tight proto-brown dwarf binary candidate. The VLA data reveals the detection of a single object closer to the western ALMA source, and with a spectral index consistent with a radio jet.
@@
Very low luminosity objects (VeLLOs), discovered by the Spitzer space telescope, are defined as embedded objects with internal luminosities () below 0.1 (di Francesco et al., 2007). They have been identified as promising proto-brown dwarf (proto-BD) candidates, that is, BDs at the earliest stages of their evolution when they are still embedded in their parental clouds (see review by Palau et al., 2024). The multiplicity of proto-BDs is an important output of substellar formation theories, and its study can help to shed light on their dominant formation mechanism (e.g. Bate, 2012; Offner et al., 2016).
We designed a pilot project to study the multiplicity of VeLLOs with ALMA, obtaining observations for IRAM 04191+1522 (IRAM04191, hereafter). IRAM04191 is a Class 0 object located in the southern part of the Taurus molecular cloud, displaying a collimated CO(2-1) bipolar outflow and an ionized jet (André et al., 1999). It was detected by the Spitzer telescope at wavelengths shorter than 60 m, and classified as a VeLLO based on its of 0.080.04 (Dunham et al., 2006). In this letter, we present unprecedented angular resolution ALMA observations of IRAM04191 at 0.89 mm that have allowed us to resolve the object into a tight binary candidate. We have complemented these data with new Very Large Array (VLA) observations, and ALMA Band 6 (1.3 mm) archival data.
@@ 2.1 Band 7 ALMA observations
IRAM04191 was observed with the ALMA 12m-array on 24 November 2017, as part of the ALMA Cycle 5 program 2017.1.00551.S (PI Huélamo). The data were obtained with 49 antennas in single field dual polarization mode. Band 7 (334.9 GHz, 0.89 mm) was used, with two basebands centered at 335.9 and 334.0 GHz, respectively, both with a width of 1.875 GHz and a channel spacing of 31 MHz. We also placed two spectral windows centered at a frequency that contains the CO(3-2) and H13CO+(4-3) transitions, respectively, to obtain serendipitous detections of gas emission associated to outflow and disk phenomena. The observations were performed with an average precipitable water-vapor column of 0.6 mm. The antenna baselines ranged from 92 m to 8.5 km, providing a maximum recoverable scale of 06. QSO J0237+2848 and QSO J0440+1437 were used as bandpass and phase calibrator, respectively. The former was used to scale the flux density. The science time on-source was 16 minutes. We used the automatic self-calibration pipeline auto_selfcal 111https://github.com/jjtobin/auto_selfcal to self-calibrate the visibilities and perform continuum imaging (Tobin et al., 2025), changing the minsnr_to_proceed parameter to be 1.5 instead of 2.95. The pipeline was executed using the Common Astronomy Software Applications package (CASA, McMullin et al., 2007) version 6.6.4. The data self-calibrated successfully, with the phase-only self-calibration converging down to the scan-length solution interval. We analyzed data with different imaging options (see Appendix B), selecting the image with robust and a cut in uvrange of 200 as our fiducial image (synthesized beam size of 3325 mas, and PA of 42°).
@@ 2.2 VLA observations
We observed IRAM04191 with the VLA of the National Radio Astronomy Observatory (NRAO) in its A configuration at K, Ka and Q bands (project 24B-250). All observations were performed during the second semester of 2024; October 22 for Q band, and December 4 and 5 for Ka and K, respectively. At all bands, we used the standard continuum observing setup which allows us to use a total of 8 GHz bandwidth for each band. For all bands, 3C147 was used as flux calibrator, 3C84 as the bandpass calibrator, and J0409+1217 as complex gain calibrator. At each band, we used cycles between target and complex gain calibrator as recommended for A configuration by the NRAO. Data were calibrated by the NRAO staff using CASA v6.6.1 and the VLA pipeline. We tried self-calibration of the data, but it did not work due to the faintness of the source. Images were made using tclean from the calibrated data and exploring different visibility weightings (natural, uniform and briggs). We have produced a single image combining all three bands (KKaQ) and using multi-frequency synthesis. We have worked with the image with a robust parameter of 1, showing a synthesized beam size of 008007, and PA of -33°.
@@ 2.3 Band 6 ALMA archival observations
We have complemented our 0.89 mm data with public ALMA archival data obtained in Band 6 (220.6 GHz, 1.3 mm) under projects 2016.1.00039.S (P.I. Dunham), and 2016.1.01284.S (P.I. Maury). We note that the former was observed with a comparable angular resolution as our 0.89 mm data. The description of these observations and the data reduction are included in Appendix A.
@@ 3.1 ALMA continuum data: most extended configurations
The 0.89 mm image of IRAM04191 is displayed in Figure 1 (left panel). The continuum emission is extended in the East-West direction, showing two bright sources separated by 54 mas (8 au at 140 pc). The strongest emission is coming from the easternmost source (IRAM04191E, hereafter). The two sources are surrounded by extended emission.
@@
The positions and fluxes for the two detected sources in the two bands are included in Table 1. Note that, in the case of the 0.89 mm data, we have derived these parameters both in the self-calibrated (Fig. 1) and non self-calibrated images (see Appendix B for details).
The analysis of the ALMA data suggests that IRAM 04191 could be part of a tight binary system. This is consistent with the fact that two ALMA sources separated by 80 mas are detected as compact sources in the two bands after applying a cut in uvrange of 1100 (see Figure 6).
We have estimated the masses (gas plus dust) of the detected sources following the procedure described in Appendix D. Both show very low values between 1.8-0.4 at 1.3 mm, and below 1 at 0.89 mm (see Table 1). The mass of the whole emission (binary plus extended emission) is estimated to be 4-5 at 0.89 mm, and 4.1 at 1.3 mm (see details in Appendix D).
@@ 3.2 VLA observations
The combined KKaQ-band image is displayed in the right panel of Figure 1. We have detected a single source at the position of IRAM04191 in the three observed bands (the three individual images are displayed in Figure 7). The properties of this radio source are included in Table 2. The image shows extended emission with an elongation towards the NE in the direction of the reported outflow (André et al., 1999). Using the fluxes at 33 and 22 GHz, we derive a spectral index of 0.240.23, consistent with free-free emission from a radio jet.
The bulk of the VLA emission seems associated to IRAM04191W, but the moderate angular resolution and sensitivity of the VLA data does not allow us to discern if the eastern component is also contributing to the centimeter emission. The position of the VLA source is offset by 30 mas south-east from IRAM04191W, which corresponds to the direction of the proper motion of Taurus members (e.g. Galli et al., 2019), although it is smaller than that reported for the neighboring moving groups (e.g. L1551, L1558), considering the time span of 7 years between the ALMA and VLA data. However, the different beam sizes and source morphologies in the ALMA and VLA data prevent a reliable determination of the proper motion. Dedicated observations are therefore required to characterize the nature of this positional offset and the proper motion of IRAM04191.
@@ 3.3 Molecular emission from C18O (2-1)
Maury (2020) has already suggested the presence of small disk around IRAM04191, so we have reanalyzed C18O (2-1) observations to shed light on the system architecture. The moment 1 is displayed in Fig. 2, where we have selected the velocity channels that show emission close to the ALMA detections: 3.5-5.5 km/s and 7.9 – 9.4 km/s. The velocity gradient in C18O is consistent with material in rotation around the ALMA binary, confirming that it may be surrounded by a circumbinary disk whose geometrical center lies closer to the faintest source (IRAM04191W). The deconvolved size of the moment 0 of the C18O (2-1) emission is 039005 0004 (PA of 888), which translates into a disk radius of 27 au, and an inclination222Estimated using the size ratio between major and minor disk axes following the expression cos-1 (/) of 647, in good agreement with the values of 50-60 derived in previous works (e.g. André et al., 1999).
We have used the moment 1 of the C18O emission to generate a diagram along its major axis which, as shown in Fig. 3 shows a Keplerian-like pattern. We have fitted the -diagram using the SLAM code (Sai, 2023), and adopting both the edge and ridge methods. We assumed a rms noise of 2.45 mJy, an SNR threshold of 5, and Gaussian weighting. The ridge method provides a more robust tracing of the emission, yielding a break radius of 11 au and a dynamical mass of .
@@ 4 Discussion
The ALMA data presented here has allowed us for the first time to (i) resolve IRAM04191 into a tight Class 0 substellar binary candidate with 11 au of projected separation surrounded by a circumbinary disk centered on the faintest ALMA continuum source (IRAM04191W), and (ii) estimate a dynamical of 5040 for the whole system. In the context of Class 0 binary systems, IRAM04191 would represent one of the tightest binaries detected in continuum emission to date, only comparable to L1157 MMS, a 16 au Class 0 binary (Tobin et al., 2022).
@@
A two-peak morphology in the continuum emission of a young object can be also interpreted as a high-inclined disk with a large inner cavity. However, such a large cavity would be hard to reconcile with the strong molecular outflow reported in this source, and the presence of a radio jet normally associated with very young objects. Moreover, in an edge-on disk scenario, the VLA centimeter detection presented here would be expected to lie between the two ALMA peaks, which is not the case.
@@
Considering that IRAM04191W seems associated to the VLA detection and its location at the center of the circumbinary disk, it is reasonable to suggest that this source may be the primary (and hence, most massive) component of the binary, despite being the fainter emitter in the ALMA data. A comparable configuration was observed in the protostellar binary [BHB2007] 11 (Alves et al., 2019), although in this case both components were detected in the VLA data. In any case, this hypothesis needs further confirmation through future dedicated ALMA and VLA observations.
@@
Palau et al. (2024) presented a correlation between the dynamical mass () of protostars and VeLLOs with their . We have represented IRAM04191 in this diagram taking advantage of the mass derived in this work, and it is found to be consistent with the established correlation, strongly suggesting its substellar nature (Fig. 4). Since the object is a binary, the individual values for both parameters should be divided taking into account the binary mass ratio, which is currently unknown.
@@
The reported values of for IRAM04191 vary between 0.05, using Herschel/SPIRE data at 250 m (Kim et al., 2016), and 0.5 using 1.3 mm data (André et al., 1999). The discrepancy between values is mainly related with the different methodology to estimate the mass, using different radius (24” vs 60”, respectively) to integrate the flux. We have reanalyzed the SPIRE/250 m data of IRAM04191 and estimated a of 0.2 integrating the flux within a 4- contour around the source, and assuming and opacity as in Kim et al. (2016). Adopting a star formation efficiency of 30% (see e.g. Palau et al., 2024), we can infer a mass reservoir of 60 available to be accreted by the binary over its evolution. Assuming a very conservative mass of 45 for each of the binary components, and that each accretes half of the mass reservoir, the system might end up as a substellar pair. We note, however, that different works have reported episodic accretion onto the source (Lee, 2007; Anderl et al., 2020). This, combined with the occurrence of mass-loss episodes (e.g. molecular outflow, radio jet), implies that the fate of the binary cannot be definitely established.
@@
Early in the formation process, turbulent fragmentation produces binaries with initial separations wider (500 au) than those formed by disk fragmentation, although dynamical evolution can quickly change these separations to lower values in 0.1 Myr (Offner et al., 2016). In that respect, core fragmentation models by Bate (2009, 2012) can produce very low-mass binaries with small separations (40 au), although the cloud properties do not resemble those of Taurus. An observational test to shed light on the formation mechanism of IRAM04191, would be to study the misalignment of the outflows and/or disks of the two components with future more sensitive ALMA observations (Offner et al., 2016). Interestingly, most BD binaries observed in the field and star forming regions are tight systems (sep 10-20 au) with mass ratio close to 1 (e.g. Duchêne and Kraus, 2013). The confirmation of IRAM04191 as a compact binary candidate at such early evolutionary stage would raise the question whether the separation distribution observed in more evolved systems could be set at very early stages of BD binary formation.

### 19 Mar 2026 v2

@@ Abstract
Context. Very low luminosity objects (VeLLOs) in nearby star-forming regions have been identified as promising proto-brown dwarf candidates, and studying their multiplicity can shed light on the dominant formation mechanism of substellar objects.
Aims. Our aim is to study the multiplicity of the VeLLO IRAM 04191+1522.
Methods. We obtained 0.89 mm ALMA observations with a very extended configuration that achieved an angular resolution of 004 (6 au at 140 pc). We complemented these data with new VLA observations and ALMA archival data at 1.3 mm.
Results. We resolved IRAM04191+1522 into a close binary candidate for the first time. The binary is detected in the ALMA continuum data with a projected separation of 80 mas (or 11 au at a distance of 140 pc). The two sources are oriented in the east-west direction, with the eastern component being brighter and more extended than the western one, which is marginally resolved. Analysis of C18O (2-1) archival data revealed gaseous material in rotation around the binary, presumably from a circumbinary disk with 27 au of radius centered on the faintest ALMA component. A fit of the position-velocity diagram allowed us to estimate a total dynamical mass for the system of 5040 . Therefore, we classify IRAM04191 as a tight proto-brown dwarf binary candidate. The VLA data revealed the detection of a single object closer to the western ALMA source with a spectral index consistent with a radio jet.
@@
Very low luminosity objects (VeLLOs), discovered by the Spitzer space telescope, are defined as embedded objects with internal luminosities () below 0.1 (di Francesco et al., 2007). They have been identified as promising proto-brown dwarf (proto-BD) candidates, that is, BDs at the earliest stages of their evolution when they are still embedded in their parental clouds (see review by Palau et al., 2024). The multiplicity of proto-BDs is an important outcome of substellar formation theories, and its study can help shed light on their dominant formation mechanism (e.g. Bate, 2012; Offner et al., 2016).
We designed a pilot project to study the multiplicity of VeLLOs with ALMA that obtained observations for IRAM 04191+1522 (IRAM04191, hereafter). IRAM04191 is a Class 0 object located in the southern part of the Taurus molecular cloud displaying a collimated CO(2-1) bipolar outflow and an ionized jet (André et al., 1999). It was detected by the Spitzer telescope at wavelengths shorter than 60 m and classified as a VeLLO based on its of 0.080.04 (Dunham et al., 2006). In this letter, we present unprecedented angular resolution ALMA observations of IRAM04191 at 0.89 mm that have allowed us to resolve the object into a tight binary candidate. We complement these data with new Very Large Array (VLA) observations and ALMA Band 6 (1.3 mm) archival data.
@@ 2.1 Band 7 ALMA observations
As part of the ALMA Cycle 5 program 2017.1.00551.S (PI Huélamo), IRAM04191 was observed with the ALMA 12m-array on November 24, 2017. The data were obtained with 49 antennas in single field dual polarization mode. Band 7 (334.9 GHz, 0.89 mm) was used, with two basebands centered at 335.9 and 334.0 GHz, respectively, both with a width of 1.875 GHz and a channel spacing of 31 MHz. We also placed two spectral windows respectively centered at the frequency that contains the CO(3-2) and H13CO+(4-3) transitions to obtain serendipitous detections of gas emission associated with outflow and disk phenomena. The observations were performed with an average precipitable water-vapor column of 0.6 mm. The antenna baselines ranged from 92 m to 8.5 km, providing a maximum recoverable scale of 06. The objects QSO J0237+2848 and QSO J0440+1437 were respectively used as a bandpass calibrator and a phase calibrator. The former was used to scale the flux density. The science time on-source was 16 minutes. We used the automatic self-calibration pipeline auto_selfcal111https://github.com/jjtobin/auto_selfcal to self-calibrate the visibilities and perform continuum imaging (Tobin et al., 2025), changing the minsnr_to_proceed parameter to be 1.5 instead of 2.95. The pipeline was executed using the Common Astronomy Software Applications package (CASA; McMullin et al., 2007) version 6.6.4. The data self-calibrated successfully, with the phase-only self-calibration converging down to the scan-length solution interval. We analyzed data with different imaging options (see Appendix B), selecting the image with a robust and a cut in range of 200 as our fiducial image showing a synthesized beam size of 3325 mas and position angle (PA) of 42°.
@@ 2.2 VLA observations
We observed IRAM04191 with the VLA of the National Radio Astronomy Observatory (NRAO) in its A configuration at the K, Ka, and Q bands (project 24B-250). All observations were performed during the second semester of 2024: October 22 for the Q band and December 4 and 5 for Ka and K, respectively. At all bands, we used the standard continuum observing setup that allows one to use a total of 8 GHz in bandwidth for each band. For all bands, 3C147 was used as a flux calibrator, while 3C84 was used as the bandpass calibrator, and J0409+1217 was used as a complex gain calibrator. At each band, we used cycles between the target and complex gain calibrator as recommended for the A configuration by the NRAO. Data were calibrated by the NRAO staff using CASA v6.6.1 and the VLA pipeline. We tried self-calibration of the data, but it did not work due to the faintness of the source. Images were made using tclean from the calibrated data and exploring different visibility weightings (natural, uniform, and briggs). We produced a single image combining all three bands (KKaQ) and using multi-frequency synthesis. We worked with the image with a robust parameter of one, showing a synthesized beam size of 008007 and a PA of -33°.
@@ 2.3 Band 6 ALMA archival observations
We complemented our 0.89 mm data with public ALMA archival data obtained in Band 6 (220.6 GHz, 1.3 mm) under projects 2016.1.00039.S (P.I. Dunham) and 2016.1.01284.S (P.I. Maury). We note that the former was observed with an angular resolution comparable to our 0.89 mm data. The description of these observations and the data reduction are included in Appendix A.
@@ 3.1 ALMA continuum data: Most extended configurations
The 0.89 mm image of IRAM04191 is displayed in Figure 1 (left panel). The continuum emission is extended in the east-west direction, showing two bright sources separated by 54 mas (8 au at 140 pc). The strongest emission comes from the easternmost source (IRAM04191E, hereafter). The two sources are surrounded by extended emission.
@@
The positions and fluxes for the two detected sources in the two bands are included in Table 1. We note that in the case of the 0.89 mm data, we derived these parameters in both the self-calibrated (Fig. 1) and non-self-calibrated images (see Appendix B for details).
The analysis of the ALMA data suggests that IRAM 04191 could be part of a tight binary system. This is consistent with the fact that two ALMA sources separated by 80 mas were detected as compact sources in the two bands after a cut in the range of 1100 was applied (see Figure 6).
We estimated the masses (gas plus dust) of the detected sources following the procedure described in Appendix D. Both show very low values: between 1.8 and 0.4 at 1.3 mm and below 1 at 0.89 mm (see Table 1). The mass of the whole emission (binary plus extended emission) is estimated to be 4-5 at 0.89 mm and 4.1 at 1.3 mm (see details in Appendix D).
@@ 3.2 VLA observations
The combined KKaQ-band image is displayed in the right panel of Figure 1. We detected a single source at the position of IRAM04191 in the three observed bands (the three individual images are displayed in Figure 7). The properties of this radio source are included in Table 2. The image shows extended emission with an elongation toward the northeast in the direction of the reported outflow (André et al., 1999). Using the fluxes at 33 and 22 GHz, we derived a spectral index of 0.240.23, consistent with free-free emission from a radio jet.
The bulk of the VLA emission seems to be associated with IRAM04191W, but the moderate angular resolution and sensitivity of the VLA data do not allow us to discern if the eastern component also contributes to the centimeter emission. The position of the VLA source is offset by 30 mas southeast from IRAM04191W, which corresponds to the direction of the proper motion of the Taurus members (e.g. Galli et al., 2019), although it is smaller than that reported for the neighboring moving groups (e.g., L1551, L1558), considering the time span of 7 years between the ALMA and VLA data. However, the different beam sizes and source morphologies in the ALMA and VLA data prevent a reliable determination of the proper motion. Dedicated observations are therefore required to characterize the nature of this positional offset and the proper motion of IRAM04191.
@@ 3.3 Molecular emission from C18O (2-1)
Maury (2020) already suggested the presence of a small disk around IRAM04191, so we reanalyzed C18O (2-1) observations to shed light on the system architecture. The moment 1 is displayed in Fig. 2, where we selected the velocity channels that show emission close to the ALMA detections: 3.5-5.5 km/s and 7.9 – 9.4 km/s. The velocity gradient in C18O is consistent with material in rotation around the ALMA binary, confirming that it may be surrounded by a circumbinary disk whose geometrical center lies closer to the faintest source (IRAM04191W). The deconvolved size of the moment 0 of the C18O (2-1) emission is 039005 0004 (PA of 888), which translates into a disk radius of 27 au and an inclination222The inclination was estimated using the size ratio between major and minor disk axes following the expression cos-1 (/). of 647, which is in good agreement with the values of 50-60derived in previous works (e.g. André et al., 1999).
We used the moment 1 of the C18O emission to generate a position-velocity () diagram along its major axis. As shown in Fig. 3, the diagram shows a Keplerian-like pattern. We fit the diagram using the SLAM code (Sai, 2023) and adopted both the edge and ridge methods. We assumed an noise of 2.45 mJy, an S/N threshold of five, and Gaussian weighting. The ridge method provides more robust tracing of the emission, yielding a break radius of 11 au and a dynamical mass of .
@@ 4 Discussion
The ALMA data presented here allowed us for the first time to (i) resolve IRAM04191 into a tight Class 0 substellar binary candidate with 11 au of projected separation surrounded by a circumbinary disk centered on the faintest ALMA continuum source (IRAM04191W) and to (ii) estimate a dynamical mass of 5040 for the whole system. In the context of Class 0 binary systems, IRAM04191 would represent one of the tightest binaries detected in continuum emission to date, only comparable to L1157 MMS (a 16 au Class 0 binary; Tobin et al., 2022).
@@
A two-peak morphology in the continuum emission of a young object can also be interpreted as a high-inclined disk with a large inner cavity. However, such a large cavity would be hard to reconcile with the strong molecular outflow reported in this source and the presence of a radio jet, which is normally associated with very young objects. Moreover, in an edge-on disk scenario, the VLA centimeter detection presented here would be expected to lie between the two ALMA peaks, and this is not the case.
@@
Considering that IRAM04191W seems to be associated with the VLA detection and its location at the center of the circumbinary disk, it is reasonable to suggest that this source may be the primary (and hence most massive) component of the binary, despite being the fainter emitter in the ALMA data. A comparable configuration was observed in the protostellar binary [BHB2007] 11 (Alves et al., 2019), although in that case both components were detected in the VLA data. Nevertheless, this hypothesis needs further confirmation through future dedicated ALMA and VLA observations.
@@
Palau et al. (2024) presented a correlation between the dynamical mass () of protostars and VeLLOs with their . Using the dynamical mass derived in this work, we included IRAM04191 in their - diagram and found it to be consistent with the reported correlation, strongly supporting its substellar nature (Fig. 4). If confirmed as a binary, the value of each parameter should be divided between components taking into account the binary mass ratio, which is currently unknown.
@@
The reported values of for IRAM04191 vary between 0.05(using Herschel/SPIRE data at 250 m, Kim et al. 2016) and 0.5 (using 1.3 mm data, André et al. 1999). The discrepancy between the values is mainly related to the methodology used to estimate the mass, which use a different radius (24” vs. 60”, respectively) to integrate the flux. We reanalyzed the SPIRE/250 m data of IRAM04191 and estimated a of 0.2 by integrating the flux within a 4 contour around the source and assuming and opacity as in Kim et al. (2016). By adopting a star formation efficiency of 30% (see, e.g., Palau et al., 2024), we can infer a mass reservoir of 60 available to be accreted by the binary over its evolution. Assuming a very conservative mass of 45 for each of the binary components and that each accretes half of the mass reservoir, the system might end up as a substellar pair. We note, however, that different works have reported episodic accretion onto the source (Lee, 2007; Anderl et al., 2020). This combined with the occurrence of mass-loss episodes (e.g., molecular outflow, radio jet) implies that the fate of the binary cannot be definitively established.
@@
Early in the stellar formation process, turbulent fragmentation produced binaries with initial separations wider (500 au) than those formed by disk fragmentation, although dynamical evolution can quickly change these separations into lower values in 0.1 Myr (Offner et al., 2016). In that respect, core fragmentation models by Bate (2009, 2012) can produce very low mass binaries with small separations (40 au), although the modelled cloud properties do not resemble those of Taurus. An observational test that could shed light on the formation mechanism of IRAM04191 would involve studying the misalignment of the outflows and/or disks of the two components with more sensitive ALMA observations in the future (Offner et al., 2016). Interestingly, most BD binaries observed in the field and star-forming regions are tight systems (sep 10-20 au) with a mass ratio close to one (e.g. Duchêne and Kraus, 2013). The confirmation of IRAM04191 as a compact binary candidate at such an early evolutionary stage would raise the question of whether the separation distribution observed in more evolved systems is already set at very early stages of BD binary formation.